OpenStudy (anonymous):

simplify: 6ln2+ln(x^2-9)-ln(x+3)

OpenStudy (anonymous):

\[\ln(\frac{2^6(x^2-9)}{x+3})\] is a start. then factor and cancel

OpenStudy (anonymous):

you get \[\ln(2^6(x-3))\]

OpenStudy (anonymous):

the instructions mean "write as a single log" not "simplify"